### v1

# references.bib

Abstract
We consider the problem of learning the dynamics in the topology of time-evolving point clouds, the prevalent spatiotemporal model for systems exhibiting collective behavior, such as swarms of insects and birds or particles in physics. In such systems, patterns emerge from (local) interactions among self-propelled entities. While several well-understood governing equations for motion and interaction exist, they are difficult to fit to data due to the often large number of entities and missing correspondences between the observation times, which may also not be equidistant. To evade such confounding factors, we investigate collective behavior from a topological perspective, but instead of summarizing entire observation sequences (as in prior work), we propose learning a latent dynamical model from topological features per time point. The latter is then used to formulate a downstream regression task to predict the parametrization of some a priori specified governing equation. We implement this idea based on a latent ODE learned from vectorized (static) persistence diagrams and show that this modeling choice is justified by a combination of recent stability results for persistent homology. Various (ablation) experiments not only demonstrate the relevance of each individual model component, but provide compelling empirical evidence that our proposed model – neural persistence dynamics – substantially outperforms the state-of-the-art across a diverse set of parameter regression tasks.
1 Introduction
Understanding emerging behavioral patterns of a collective through the interaction of individual entities is key to elucidate many phenomena in nature on a macroscopic and microscopic scale. Prominent examples are coherently moving flocks of birds, the swarming behavior of insects and fish or the development of cancerous cells, all understood as 2D/3D point clouds that evolve over time. Importantly, several widely-accepted governing equations for collective behavior exist [Dorsogna06a, Vicsek95a, Motsch18a] which, when appropriately parameterized, can reproduce different incarnations of typically observed patterns. Even more importantly, these equations are tied to physically interpretable parameters and can provide detailed insights into the intrinsic mechanisms that control various behavioral regimes. However, while it is fairly straightforward to simulate collective behavior from governing equations (see [Diez21a]), the inverse problem, i.e., identifying the model parameters from the data, turns out to be inherently difficult. Confounding factors include the often large number of observed entities and the difficulty of reliably identifying individual trajectories due to missing correspondences between individuals at possibly non-equidistant observation times.
Figure 1: Neural Persistence Dynamics. Given is a sequence of observed point clouds $\mathcal{P}_{\tau_{0}},\ldots,\mathcal{P}_{\tau_{N}}$. First, we summarize each $\mathcal{P}_{\tau_{i}}$ via (Vietoris-Rips) persistent homology into persistence diagrams (zero-, one- and two-dimensional; only the zero-dimensional diagrams are shown here) which are then vectorized into $\mathbf{v}_{\tau_{i}}$ via existing techniques. Second, we model the dynamics in the sequence $\mathbf{v}_{\tau_{0}},\ldots,\mathbf{v}_{\tau_{N}}$ via a continuous latent variable model (in our case, a latent ODE) and then use the latent path to predict the parameters of well-established governing equation(s) of collective behavior. Precomputed steps are highlighted in red (best-viewed in color).
However, as several works [Topaz15a, Bhaskar19a, Giusti23a] have recently demonstrated, it may not be necessary to rely on individual trajectories for parameter identification. In fact, collective behavior is characterized by global patterns that emerge from local interactions, and we observe the emergence of these patterns through changes to the “shape” of point clouds over time. For instance, birds may form a flock, split into groups, and then merge again. This perspective has prompted the idea of summarizing such topological events over time and then phrasing model identification as a downstream prediction/regression task. The key challenge here lies in the transition from topological summaries of point clouds at specific observation times, typically obtained via persistent homology (PH) [Barannikov94a, Carlsson21a, Edelsbrunner2010Computational], to the dynamic regime where the temporal dimension plays a crucial role.
[…]
vicsek-10k
Ours (PointNet++, v2) 0.272$\pm$0.083 0.199$\pm$0.014
Ours (PH, v1) 0.576$\pm$0.032 0.146$\pm$0.006
dorsogna-1k
Ours (PointNet++, v2) 0.814$\pm$0.032 0.132$\pm$0.018
Ours (PH, v1) 0.846$\pm$0.011 0.097$\pm$0.005
Table 1: PointNet++ [Qi17a] vs. persistent homology (PH) representations; $\uparrow$ means higher and $\downarrow$ means lower is better.
Yet, despite the often remarkable performance of topological approaches in terms of parameter identification for models of collective behavior, it remains unclear in which situations they are preferable over more traditional (learning) methods that can handle point cloud data, as, e.g., used in computer vision problems [Rempe20a, Guo20a]. In the spirit of [Turkes22a], we highlight this point by previewing a snapshot of one ablation experiment from Sec. 4. In particular, Tbl. 1 compares the parameter regression performance of our proposed approach, using PH, to a variant where we instead use (jointly learned) PointNet++ [Qi17a] representations. Referring to Fig. 1, this means that the $\mathbf{v}_{\tau_{i}}$ are computed by a PointNet++ model. As can be seen in Tbl. 1, relying on representations computed via PH works well on both datasets, while PointNet++ representations fail on one dataset (vicsek-10k), but indeed perform well on another (dorsogna-1k).
In view of this observation, it is worth highlighting, that there is a clear distinction in terms of the source of point cloud dynamics when comparing collective behavior to problems in computer vision where moving objects are of primary interest and PointNet variants (see [Guo20a]) have shown stellar performance: in particular, point clouds in vision primarily evolve due to a change in pose or camera motion, whereas collective behavior is driven by (local) intrinsic point interactions. The latter induces changes to the “shape” of the point clouds, i.e., a phenomenon that is typically not observed in vision.
Contribution(s). Our key idea is to learn a generative model that can reproduce topological summaries at each point in time. This contrasts prior work, which primarily aims to extract one summary representation of the entire timeline. In detail, we advocate for modeling the dynamics of vectorized persistence diagrams via a continuous latent variable model (e.g., a latent ODE), see Fig. 1, and to use the resulting latent paths as input to a downstream parameter regression task. Recent stability results for vectorizations of persistence diagrams – relating distances among the latter to the Wasserstein distance between point clouds – justify this modeling choice. Aside from state-of-the-art performance on various parameter identification tasks for established models of collective behavior, our approach scales favorably with the number of observed sequences, accounts for non-equidistant observation times, and is easily combinable with other sources of information.
2 Related work
Our work is partially related to the literature on learning with dynamic point clouds in vision problems, but primarily connects to work on summarizing topological features over time and inferring interaction laws of collective dynamics from data.
Summarizing topological features over time. A common denominator in the literature on encoding topological changes in dynamic point clouds is the use of persistent homology, extended to accommodate the temporal dimension in various ways. One of the early works along this direction are persistence vineyards [CohenSteiner06a], introduced as a means to study folding trajectories of proteins. Based on the idea of tracking points in persistence diagrams over time, vineyards come with stability properties similar to persistent homology [Munch13a], but are expensive to compute and compare on a large scale. In cases where one would know the “right” scale at which to compute homology, one may also use zigzag persistence [Carlsson10a], as done in [Corcoran17a, Tymochko20a], to track homological changes over time. Yet, for problems with many observation sequences, scale selection per sequence is nontrivial and highly impractical.
Alternatively, instead of thinking about the evolution of individual points in persistence diagrams one may discard the matching between points and instead consider sequences of summary representations of said diagrams. In [Hajij18a], the authors work directly with persistence diagrams to identify changes in the topology of time-varying graphs. In terms of temporal summary representations, [Topaz15a] introduce crocker plots to encode the evolution of topological features by stacking discretized Betti curves over time. Crocker stacks [Xian22a], an extension of this concept, adds a smoothing step that gradually reduces the impact of points of low persistence and, upon discretization, yields a third dimension to crocker plots. Both, crocker plots & stacks have been used as input to regression methods to predict the configuration of models of collective behavior [Bhaskar19a]. By drawing on prior work on kernels for sequentially ordered data [Kiraly19a][Giusti23a] follow a conceptually similar strategy as [Topaz15a, Xian22a], introducing a path signature kernel (PSK) for sequences of summary representations of persistence diagrams.
Along a different line of research, [Kim18a, Kim18b] propose formigrams as summaries of dynamic metric data, encoding the evolution of connected components. In subsequent work, [Kim21a] construct multidimensional (i.e., spatiotemporal) persistent homology modules from dynamic metric data and compare invariants of these modules. While these works provide important theoretical stability results of dimension zero, i.e., connected components, it is unclear how their construction extends to homological features of higher dimension in a tractable manner. However, it is worth pointing out that recent progress for vectorizations of multiparameter persistent homology [Loiseaux2023a] in combination with [Kim21a] (who essentially construct multiparameter persistence modules) might, in the future, constitute a tractable approach to study dynamically changing point clouds with learning methods.
Notably, computational challenges also arise in the context of crocker plots & stacks and PSKs. Despite remarkable success in distinguishing different configurations of models for collective behavior, all approaches suffer scalability issues, either in terms of the dimensionality of the vectorized persistence diagrams (as with the PSK approach of [Giusti23a]), or in terms of the number of observation sequences (as is the case for crocker plots/stacks, due to the need for extensive cross-validation of the discretization parameters). As we show in Sec. 4, our method outperforms these techniques by a large margin, scales to large amounts of training sequences and requires little hyperparameter tuning.
Finally, aside from the closely-related works discussed above, we highlight that there is a large body of literature on the topological analysis of time-varying signals, such as studying fMRI data via cubical persistence [Rieck20a], or the persistent homology of delay embeddings [Perea15a, Perea15b] of time series. In our context, these works are only partially related, though, as they assume knowledge about individual trajectories over time (e.g., voxel IDs in fMRI data), which is unrealistic in our experimental setting.
Inferring interaction laws for models of collective behavior. In the context of capturing the characteristics of collective behavior, there is a second line of closely related work on inferring and studying interaction laws (see, e.g., [Brunton16a, Ballerin08a, Bialek12a, Katz10a, Lu19a, Zhong20a]), ranging from metric-distance-based models and topological interaction models, to non-parametric estimators of interaction kernels. While a full survey of this literature is beyond the scope of this paper, we highlight that one common denominator in these works is their reliance on correspondences between points across time, e.g., to infer individual point velocities or trajectories. To give an example, in recent work [Zhong20a, Lu19a], the authors derive non-parametric estimators for interaction kernels (certain functions of pairwise distances) between observed points which crucially hinges on the traceability of each particle over time. While such approaches are conceptually appealing and even allow for reconstructing or continuing trajectories, they operate in the observation space. Even under a moderate number of points $m$, computing these estimators becomes prohibitively expensive (as, e.g., in 3D, the state space is $\mathbb{R}^{3m}$). In contrast, our approach only requires positional information, and can handle much larger point clouds. Furthermore, although we only present results on predicting parameters for a class of a priori given governing equations, by formulating an auxiliary regression task, our underlying model may also be used to predict other quantities of interest.
3 Method
Below, we present our method with a focus on its application to modeling collective behavior in point clouds. Note, however, that the latter represents only one of many conceivable use cases, as our method can be applied in exactly the same way, whenever one can extract topological features from time-dependent data (e.g., graphs or images).
Notation. In the following, $\mathcal{P}=\{\mathbf{x}_{1},\ldots,\mathbf{x}_{M}\}\subset\mathbb{R}^{3}$ denotes a point cloud with $M$ points and $d(\mathbf{x},\mathbf{y})=\|\mathbf{x}-\mathbf{y}\|$ denotes the Euclidean metric. Point clouds may be indexed by $t_{i}$ (or $\tau_{i}$) to highlight the dependence on time $t_{i}$. If necessary, we clearly distinguish between $\tau_{i}$ as a time point with an available observation, and $t_{i}$ as a general placeholder for time.
Problem statement. Given a sequence of point clouds $\mathcal{P}_{\tau_{0}},\ldots,\mathcal{P}_{\tau_{N}}$, observed at possibly non-equidistant time points $\tau_{i}$, we (1) seek to model their topological evolution over time and then (2) use this model to predict the parametrization of an a priori defined governing equation of collective behavior. The latter is typically specified by a small set of parameters $(\beta_{1},\ldots,\beta_{P})$ that control the motions $\mathrm{d}\mathbf{x}_{i}/\mathrm{d}t$ of individual points $\mathbf{x}_{i}$ and specify (local) interactions among neighboring points.
As a preliminary introduction to our model description, we briefly establish how one may extract topological features from a point cloud $\mathcal{P}$, using persistent homology [Edelsbrunner2010Computational, Barannikov94a, Carlsson21a] – the arguably most prominent and computationally most feasible approach.
Persistent homology of point clouds. Persistent homology seeks to uncover and concisely summarize topological features of $\mathcal{P}$. To this end, one constructs a topological space from $\mathcal{P}$ in the form of a simplicial complex, and studies its homology across multiple scales. The most relevant construction for our purposes is the Vietoris-Rips complex $\operatorname{Rips}(\mathcal{P})_{\delta}$, with vertex set $\mathcal{P}$. This complex includes an $m$-simplex $[\mathbf{x}_{0},\ldots,\mathbf{x}_{m}]$ iff $d(\mathbf{x}_{i},\mathbf{x}_{j})\leq\delta$ for all $0\leq i,j\leq m$ at a given threshold $\delta$. The “shape” of this complex can then be studied using homology, a tool from algebraic topology, with zero-dimensional homology ($H_{0}$) encoding information about connected components, one-dimensional homology ($H_{1}$) encoding information about loops and two-dimensional homology ($H_{2}$) encoding information about voids; we refer to this information as homological/topological features. Importantly, if $\delta_{b}\leq\delta_{d}$, then $\operatorname{Rips}(\mathcal{P})_{\delta_{b}}\subseteq\operatorname{Rips}(\mathcal{P})_{\delta_{d}}$, inducing a sequence of inclusions when varying $\delta$, called a filtration. The latter in turn induces a sequence of Abelian groups (or vector spaces) $\operatorname{H}_{k}(\operatorname{Rips}(\mathcal{P})_{\delta_{b}})\to\operatorname{H}_{k}(\operatorname{Rips}(\mathcal{P})_{\delta_{d}})$ at the homology level. Throughout this sequence, homological features (the generators of the groups) may appear and disappear, we say they are born at some $\delta_{b}$ and die at $\delta_{d}>\delta_{b}$. For instance, a one-dimensional hole might appear at a particular value of $\delta_{b}$ and disappear at a later $\delta_{d}$; in other words, it persists from $\delta_{b}$ to $\delta_{d}$, hence the name persistent homology. As different features may be born and die at the same time, the collection of (birth, death) tuples is a multiset of points, often represented in the form of a persistence diagram, which we denote as $\operatorname{dgm}_{k}(\operatorname{Rips}(\mathcal{P}))$. Here, the notation $\operatorname{Rips}(\mathcal{P})$ refers to the full filtration and $\operatorname{dgm}_{k}$ indicates that we have tracked $k$-dimensional homological features throughout this filtration.
Clearly this construction does not account for any temporal changes to a point cloud, but reveals features that are present at a specific time point. Furthermore, due to the inconvenient multiset structure of persistence diagrams for learning problems, one typically resorts to appropriate vectorizations (see, e.g., [Bubenik15a, Adams17a, Hofer19b, Carriere20a]), all accounting for the fact that points close to the diagonal $\Delta=\{(x,x):x\in\mathbb{R}\}$ contribute less (0 at $\Delta$) to the vectorization. The latter is important to preserve stability (see paragraph below) with respect to perturbations of points in the diagram. In the following, we refer to a vectorized persistence diagram of a point cloud $\mathcal{P}_{t_{i}}$ as
$\mathbf{v}_{t_{i},k}:=\operatorname{vec}(\operatorname{dgm}_{k}(\operatorname{Rips}(\mathcal{P}_{t_{i}})))\enspace,$ (1)
where $\mathbf{v}_{t_{i},k}\in\mathbb{R}^{d}$ and $d$ is controlled by hyperparameter(s) of the chosen vectorization technique. For brevity, we omit the subscript $k$ when referring to vectorization unless necessary.
Remark 1.
As our goal is to model the dynamics of vectorized persistence diagrams using a continuous latent variable model, it is important to discuss the dependence of the vectorizations on the input data. In particular, for our modeling choice to be sound, vectorizations should vary (Lipschitz) continuously with changes in the point clouds over time. We discuss this aspect next.
Stability/Continuity aspects. First, we point out that persistence diagrams can be equipped with different notions of distance, most prominently the bottleneck distance and Wasserstein distances, both based upon the cost of an optimal matching between two diagrams, allowing for matches to the diagonal $\Delta$. We refer the reader to [Chazal14d, Skraba23a] for a detailed review. Stability in the context of persistent homology is understood as persistence diagrams varying (Lipschitz) continuously with the input data. While seminal stability results exist for the Wasserstein distances [CohenSteiner10a] and the bottleneck distance [CohenSteiner07a, Chazal14d], most results from the literature focus on the latter.
In case of vectorization techniques for persistence diagrams (e.g., [Adams17a, Hofer19b]), it turns out that, in most cases, vectorizations are only Lipschitz continuous with respect to Wasserstein distances, and existing results are typically of the form
$d\!\left(\operatorname{vec}(F),\operatorname{vec}(G)\right)\leq K\,W_{1}(F,G)\enspace,$ (2)
where $W_{1}$ denotes the $(1,2)$-Wasserstein distance as defined in [Skraba23a, Def. 2.7], $F,G$ are two persistence diagrams and $K>0$ is a Lipschitz constant. In the context of Rem. 1, one may be tempted to combine Eq. 2 with the seminal Wasserstein stability theorem from [CohenSteiner10a] to infer stability of vectorizations with respect to the input data. Yet, the conditions imposed in [CohenSteiner10a] actually eliminate a direct application of this result in most practical cases, as discussed in [Skraba23a]. However, the latter work also provides an alternative: in our particular case of Vietoris-Rips persistent homology, one can upper bound $W_{1}$ in terms of the standard point set Wasserstein distance $\mathcal{W}_{1}$. Specifically, upon relying on, e.g., the vectorization technique from [Hofer19b], we obtain the inequality chain
[…]
where we set $\operatorname{vec}\equiv s_{\boldsymbol{\lambda}}$ as in [Hofer19b, Def. 9], and $F_{k}=\operatorname{dgm}_{k}(\operatorname{Rips}(\mathcal{P}))$, $G_{k}=\operatorname{dgm}_{k}(\operatorname{Rips}(\mathcal{Q}))$ denote the $k$-dimensional persistence diagrams of point clouds $\mathcal{P}$ and $\mathcal{Q}$ of equal cardinality $M$.
This continuity property guarantees that small changes in dynamic point clouds over time only induce small changes in their vectorized persistence diagrams, and therefore provides a solid justification for the model discussed next.
Latent variable model for persistence vectorizations. As, we presume that the dynamic point clouds under consideration are produced (or can be described sufficiently well) by equations of motions with only a few parameters, cf. Fig. 3, it is reasonable to assume that the dynamics of the (vectorized) persistence diagrams are equally governed by a somewhat simpler unobserved/latent dynamic process in $\mathbb{R}^{z}$ with $z\ll d$. We model this latent dynamic process via a neural ODE [Chen18a, Rubanova19a], learned in a variational Bayes regime [^1] Other choices, e.g., a latent SDE as in [Li20a, Zeng23a] are possible, but we did not explore this here. [Kingma14a]. In this setting, one chooses a recognition/encoder network (Encθ) to parametrize an approximate variational posterior $q_{\boldsymbol{\theta}}(\mathbf{z}_{t_{0}}|\{\mathbf{v}_{\tau_{i}}\}_{i})$, an ODE solver to yield $\{\mathbf{z}_{\tau_{i}}\}_{i}$ at observed time points $\tau_{i}\in[0,T]$ and a suitable generative/decoder network (Decγ) to implement the likelihood $p_{\boldsymbol{\gamma}}(\mathbf{v}_{\tau_{i}}|\mathbf{z}_{\tau_{i}})$. Upon choosing a suitable prior $p(\mathbf{z}_{t_{0}})$, one can then train the model via ELBO maximization, i.e.,
$\boldsymbol{\theta},\boldsymbol{\gamma}=\arg\max_{\boldsymbol{\theta},\boldsymbol{\gamma}}\mathbb{E}_{\mathbf{z}_{t_{0}}\vspace{-0.2cm}\sim q_{\boldsymbol{\theta}}}\Big{[}\sum\mathop{}_{\mkern-5.0mui}\log p_{\boldsymbol{\gamma}}(\mathbf{v}_{\tau_{i}}|\mathbf{z}_{\tau_{i}})\Big{]}-\textit{D}_{\text{KL}}(q_{\boldsymbol{\theta}}(\mathbf{z}_{t_{0}}|\{\mathbf{v}_{\tau_{i}}\}_{i})\|\,p(\mathbf{z}_{t_{0}}))\enspace.$ (4)
Different to [Rubanova19a], we do not implement the recognition/encoder network via another neural ODE, but rather choose an attention-based approach (mTAN) [Shukla21a] which has proven to be very effective in practice [Zeng23a], and can even be used in a standalone manner as a strong baseline (see Sec. 4). In our implementation, the recognition network yields the parametrization ($\boldsymbol{\mu},\boldsymbol{\Sigma}$) of a multivariate Gaussian in $\mathbb{R}^{z}$ with diagonal covariance, and the prior is a standard Gaussian $\mathcal{N}(\mathbf{0},\mathbf{I}_{z})$. Furthermore, the ODE solver (e.g., Euler) can yield $\mathbf{z}_{t_{i}}$ at any desired $t_{i}$, however, we can only evaluate the ELBO at observed $\tau_{i}$.
Figure 2: Schematic illustration of different model variants. The first three variants (left to right) explicitly model latent dynamics (later denoted as w/ dynamics), the baseline variants do not (later denoted as w/o dynamics), but still incorporate the attention mechanism of the encoder from [Shukla21a], which we use throughout.
Regression objective. To realize the downstream regression task, i.e., regressing parameters of an underlying governing equation for collective behavior, we have multiple choices. By our assumption of a latent dynamic process that carries information about the dynamic nature of the topological changes, it is reasonable to tie the regression objective to the latent path. In particular, we can even re-use the attention-based encoder architecture to allow attending to different parts of this path. However, different to $\text{{Enc}}_{\boldsymbol{\theta}}$, which parametrizes the approximate posterior from observations at time points $\tau_{i}$, the regression network $\text{{Enc}}_{\boldsymbol{\alpha}}$ (with its own set of parameters $\boldsymbol{\alpha}$) receives latent states at equidistant $t_{i}\in[0,T]$. As in [Rubanova19a], for training, we extend the ELBO objective of Eq. 4 by an additive auxiliary regression loss, such as the mean-squared error.
A schematic overview of our neural persistence dynamics approach is shown in Fig. 1; additional details, including different model variants, are illustrated in Fig. 2.
4 Experiments
4.1 Datasets
Similar to previous work [Topaz15a, Bhaskar19a, Xian22a, Giusti23a], we simulate models of collective behavior under different parametrizations to produce a large corpus of observation sequences with varying dynamics and varying emergent behavior. For simulation, we use SiSyPHE [Diez21a] [^2] https://github.com/antoinediez/Sisyphe and consider three different models in $\mathbb{R}^{3}$: D’Orsogna [Dorsogna06a], Vicsek [Degond08a] and volume exclusion [Motsch18a]. The corresponding equations of motion (and interaction laws) are summarized in Fig. 3 with parameters that are varied in our experiments highlighted in red. We sample these parameters, as specified in Appendix A, so that a wide range of macroscopic behavioral regimes is covered. For each sampled parameter configuration, we simulate one sequence of point clouds, to a total of 10,000 sequences per model. All simulations are run for 1,000 time steps (with step size 0.01, starting at $t=0$) on point clouds of size $M=200$. We take every 10th time step as an observation, yielding observation sequences of length 100. At $t=0$, points are drawn uniformly in $[-0.5,0.5]^{3}$, and initial velocities are uniformly distributed on the unit sphere. For direct comparison to [Giusti23a], we also simulated their parameter configuration of the D’Orsogna model. This setup yields four datasets: dorsogna-1k (from [Giusti23a]), dorsogna-10k, vicsek-10k and volex-10k.
D’Orsogna (dorsogna) [Dorsogna06a] $\frac{\mathrm{d}\mathbf{x}_{i}}{\mathrm{d}t}\,{=}\,\mathbf{v}_{i},\,{\color[rgb]{0.83984375,0.15234375,0.15625}\definecolor[named]{pgfstrokecolor}{rgb}{0.83984375,0.15234375,0.15625}m}\frac{\mathrm{d}\mathbf{v}_{i}}{{\mathrm{d}t}}\,{=}\left({\color[rgb]{0.83984375,0.15234375,0.15625}\definecolor[named]{pgfstrokecolor}{rgb}{0.83984375,0.15234375,0.15625}\alpha}{-}\beta|\mathbf{v}_{i}|^{2}\right)\mathbf{v}_{i}{-}\frac{1}{M}\nabla_{\mathbf{x}_{i}}\sum_{j=1,j\neq i}^{M}U(\|\mathbf{x}_{i}{-}\mathbf{x}_{j}\|),\,U(r)\,{=}\,{\color[rgb]{0.83984375,0.15234375,0.15625}\definecolor[named]{pgfstrokecolor}{rgb}{0.83984375,0.15234375,0.15625}C_{r}}e^{-\frac{r}{{\color[rgb]{0.83984375,0.15234375,0.15625}\definecolor[named]{pgfstrokecolor}{rgb}{0.83984375,0.15234375,0.15625}l_{r}}}}{-}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}C_{a}}e^{-\frac{r}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}l_{a}}}}$ (5) Vicsek (vicsek) [Degond08a, Vicsek95a] $\frac{\mathrm{d}\mathbf{x}_{i}}{\mathrm{d}t}\,{=}\,{\color[rgb]{0.83984375,0.15234375,0.15625}\definecolor[named]{pgfstrokecolor}{rgb}{0.83984375,0.15234375,0.15625}c}\,\mathbf{v}_{i},\ \mathrm{d}\mathbf{v}_{i}\,{=}\left(\mathbf{I}{-}\mathbf{v}_{i}\mathbf{v}_{i}^{\top}\right)\left({\color[rgb]{0.83984375,0.15234375,0.15625}\definecolor[named]{pgfstrokecolor}{rgb}{0.83984375,0.15234375,0.15625}\nu}\,\bar{\mathbf{v}}_{i}\,\mathrm{d}t{+}\sqrt{2{\color[rgb]{0.83984375,0.15234375,0.15625}\definecolor[named]{pgfstrokecolor}{rgb}{0.83984375,0.15234375,0.15625}D}}\,\mathrm{d}\mathbf{B}_{t}\right),\ \bar{\mathbf{v}}_{i}\,{=}\,\frac{\mathbf{a}_{i}}{\|\mathbf{a}_{i}\|},\ \mathbf{a}_{i}\,{=}\,\sum_{\hskip 5.69046ptj,\|\mathbf{x}_{i}{-}\mathbf{x}_{j}\|\leq{\color[rgb]{0.83984375,0.15234375,0.15625}\definecolor[named]{pgfstrokecolor}{rgb}{0.83984375,0.15234375,0.15625}R}}\mathbf{v}_{j}$ (6) Volume exclusion [Motsch18a] (volex) variant from [Diez21a] $\begin{split}\frac{\mathrm{d}\mathbf{x}_{i}}{\mathrm{d}t}&\,{=}\,{-}\frac{{\color[rgb]{0.83984375,0.15234375,0.15625}\definecolor[named]{pgfstrokecolor}{rgb}{0.83984375,0.15234375,0.15625}\alpha}}{{\color[rgb]{0.83984375,0.15234375,0.15625}\definecolor[named]{pgfstrokecolor}{rgb}{0.83984375,0.15234375,0.15625}R}}\sum_{j=1,i\neq j}^{M}\phi\left(\frac{\|\mathbf{x}_{j}{-}\mathbf{x}_{i}\|^{2}}{4{\color[rgb]{0.83984375,0.15234375,0.15625}\definecolor[named]{pgfstrokecolor}{rgb}{0.83984375,0.15234375,0.15625}R}^{2}}\right)(\mathbf{x}_{i}{-}\mathbf{x}_{j}),~{}~{}\text{with}~{}~{}\phi(r)=\begin{cases}\frac{1}{r}\,{-}\,1,&0<r\leq 1\\ 0,&\text{else}\end{cases}\\ &+\text{\emph{cell division} \& \emph{cell death} at constant rate(s) ${\color[rgb]{0.83984375,0.15234375,0.15625}\definecolor[named]{pgfstrokecolor}{rgb}{0.83984375,0.15234375,0.15625}\lambda_{b}},{\color[rgb]{0.83984375,0.15234375,0.15625}\definecolor[named]{pgfstrokecolor}{rgb}{0.83984375,0.15234375,0.15625}\lambda_{d}}\in[0,1]$}\end{split}$ (7)
Figure 3: Models of collective behavior. Parameters that are varied to obtain different behavior are highlighted in red; the range of each parameter is listed in Appendix A. In the Vicsek model, $\mathbf{B}_{t}$ denotes Brownian motion.
4.2 Implementation, training & evaluation metrics
In general, we randomly partition each dataset into five 80/20 splits for training and testing. On each training split, we train three models using only a fraction (i.e., 20%, 50% and 80%) of randomly chosen time points per sequence. Similarly, all testing splits undergo the same sampling procedure. Hence, our evaluation metrics (see below) provide a robust estimate of performance for different regimes of missing and unevenly spaced observations. All models are trained using ADAM [Kingma15a] with a weight decay and KL divergence weight of 0.001 for 150 epochs at a learning rate of 0.001. To compute Vietoris-Rips persistence diagrams for zero, one, and two-dimensional persistent homology, we rely on Ripser++ [Zhang20a]. Vectorizations for each diagram dimension are computed using exponential structure elements from [Hofer19b, Sec. 5.4.1] (i.e., the $s_{\boldsymbol{\lambda}}$ in Eq. 3), located at 20 $k$-means++ cluster centers and with variance(s) chosen according to [Royer21a]. Upon concatenation of the vectorizations, this yields $d=60$ dimensional representations per point cloud and time point. Latent space dimensionality $z$ is fixed to 20 and the ODE integration time is always scaled to $[0,1]$. While other settings are certainly possible, we did not observe any noticeable benefits from increasing the dimensionality of the vectorizations or the latent space.
Evaluation metrics. We report the coefficient of determination $R^{2}$ (in $[0,1]$; higher is better $\uparrow$) and the symmetric mean absolute percentage error (SMAPE, in $[0,1]$; lower is better $\downarrow$). In detail, we first collect both scores on the testing portion of our five random 80/20 splits, and all three time point sampling percentages. In each table, we then list the average of these scores with one standard deviation. Further, when reporting results on a dataset, we mark the best score in bold, as well as all other scores that do not show a statistically significant difference in mean (assessed via a Mann-Whitney test at 5% significance, and correcting for multiple comparisons).
4.3 Ablation
In our ablation study, we assess (1) the relevance of different point cloud representations (PH vs. PointNet++), (2) any potential benefits of modeling latent dynamics and (3) the impact of varying the observation timeframe.
Are representations complementary? We first investigate the impact of different point cloud representations (from PH and PointNet++, resp.), by comparing variants v1,v2 and v3 from Fig. 2.
Source $\oslash\ R^{2}\uparrow$oslash SMAPE $\downarrow$
vicsek-10k
PointNet++ (v2) 0.272$\pm$0.083 0.199$\pm$0.014
PH (v1) 0.576$\pm$0.032 0.146$\pm$0.006
PH+PointNet++ (v3) 0.575$\pm$0.030 0.144$\pm$0.006
dorsogna-1k
PointNet++ (v2) 0.814$\pm$0.032 0.132$\pm$0.018
PH (v1) 0.846$\pm$0.011 0.097$\pm$0.005
PH+PointNet++ (v3) 0.930$\pm$0.003 0.067$\pm$0.002
Table 2: Ablation study on the relevance of different point cloud representations.
Tbl. 2 presents an extension of Tbl. 1, listing results for the vicsek-10k and dorsogna-1k data. While using PointNet++ representations on vicsek-10k yields rather poor performance, combining them with representations computed from PH is at least not detrimental. On the other hand, on dorsogna-1k, where PointNet++ yields decent SMAPE and $R^{2}$ scores, the combination of both sources substantially outperforms each single source in isolation. Although, the complementary nature of a topological perspective on data has been pointed out many times in the literature, it is rarely as pronounced as in this particular experiment. Hence, we will stick to the combination of PH+PointNet++ representations (i.e., v3 in Fig. 2) in any subsequent experiments.
Are explicit latent dynamics beneficial? To assess the impact of modeling the dynamics of persistence diagram vectorizations, we ablate the latent ODE part of our model. In detail, we compare v3 from Fig. 2 against using $\texttt{Enc}_{\boldsymbol{\alpha}}$ operating directly on concatenated point cloud representations from PH and PointNet++ (i.e., a combination of v4 & v5 from Fig. 2). The latter approach constitutes an already strong baseline (cf. [Zeng23a]) as $\texttt{Enc}_{\boldsymbol{\alpha}}$ incorporates an attention mechanism that allows attending to relevant parts of each sequence. For a fair comparison, we increase the size of $\texttt{Enc}_{\boldsymbol{\alpha}}$ to approximately match the number of parameters to our latent dynamic model.
$\oslash\ R^{2}\uparrow$oslash SMAPE  $\downarrow$
vicsek-10k
w/ dynamics 0.574$\pm$0.030 0.144$\pm$0.006
w/o dynamics 0.511$\pm$0.020 0.155$\pm$0.004
dorsogna-1k
w/ dynamics 0.930$\pm$0.003 0.067$\pm$0.002
w/o dynamics 0.850$\pm$0.006 0.098$\pm$0.003
Table 3: Results on the relevance of latent dynamics.
Similar to Tbl. 2, Tbl. 3 lists results for the vicsek-10k and dorsogna-1k data. As we can see, explicitly modeling the latent dynamics is beneficial, with considerable improvements in the reported scores across both datasets. While differences are quite prominent on the dorsogna-1k data, they are less pronounced in the SMAPE score on vicsek-10k, but still statistically significant. However, it is important to point out, that even without any explicit latent dynamics, the regression performance is already above the current state-of-the-art (cf. PSK [Giusti23a] results in Tbl. 4), highlighting the necessity for strong baselines.
Impact of the observation timeframe. So far, we have presented results (as in prior work) where observation sequences are extracted from the beginning of a simulation (i.e., from $t_{0}=0$), which corresponds to observing the emergence of patterns out of a random, yet qualitatively similar, initial configuration of points. In this experiment, we investigate the impact of extracting observation sequences with starting points randomly spread across a much longer simulation time. This is relevant as it is typically unlikely to observe a random configuration of points in practice initially. To this end, we create variations of the dorsogna-10k data by progressively increasing the simulation time $T$ from 1,000 to 20,000, and then randomly extracting sub-sequences of length 1,000 (again, as before, taking each 10th step as an observation). The extension of the simulation timeframe has the effect that the difficulty of the learning task considerably increases with $T$. We argue that this is due to the increased variation across the observed sequences, while the amount of training data remains the same. For comparison, we also list results for the PSK approach of [Giusti23a], however, only for the optimal case of observing all time points within a sequence (due to the massive PSK computation time; >3days). As the PSK relies on persistent homology only, we use our v1 variant from Fig. 2. Fig. 4 displays an initial increase/drop in SMAPE and $R^{2}$, resp., (cf. Tbl. 4) for both approaches, yet our approach degrades much slower with $T$.
2,0005,00010,00020,$0$.40.40.40.40.50.50.50.50.60.60.60.6Simulation time $T$$R^{2}~{}(\uparrow)$Ours, v1PSK [Giusti23a]
2,0005,00010,00020,$0$.110.110.110.110.120.120.120.120.130.130.130.130.140.140.140.14Simulation time $T$SMAPE ()($\downarrow$)( ↓ )Ours, v1PSK [Giusti23a]
Figure 4: Impact of the maximal simulation time $T$ for extracting training/testing sequences starting at $\tau_{0}\in[0,T-1000]$.
4.4 Comparison to the state-of-the-art
Finally, we present parameter regression results on the full battery of datasets and compare against two of the most common approaches from the literature: the path signature kernel (PSK) of [Giusti23a] and the crocker stacks approach from [Xian22a]. For both approaches, we report results when all time points are observed to establish a baseline of what could be achieved in the best case. Note that we evaluate the PSK approach [Giusti23a] on exactly the same vectorizations as our approach and we replicate their experimental setting of choosing the best hyperparameters via cross-validation. Similarly, for crocker stacks, we cross-validate the discretization parameters, see Secs. B.1 and B.2. Tbl. 4 lists the corresponding results. Aside from the observation that our approach largely outperforms the state-of-the-art on these parameter regression tasks, we also highlight that our model is trained with exactly the same hyperparameters across all datasets.
Table 4: Comparison to the state-of-the-art across four diverse datasets of collective behavior. Ours (joint) refers to the variant v3 of Fig. 2, Ours (PH-only) refers to variant v1. The latter setting is directly comparable to the PSK and crocker stacks approach. Best results are marked bold. Multiple bold entries indicate that there is no significant difference in mean. Note that dorsogna-1k is simulated exactly as in [Giusti23a] varying only two parameters, whereas for dorsogna-10k we vary four parameters (as with vicsek-10k and volex-10k).
$\oslash\,R^{2}\uparrow$oslash SMAPE $\downarrow$
dorsogna-1k Ours (joint, v3) 0.930$\pm$0.003 0.067$\pm$0.002
Ours (PH-only, v1) 0.846$\pm$0.011 0.097$\pm$0.005
PSK [Giusti23a] 0.816$\pm$0.015 0.096$\pm$0.006
Crocker Stacks [Xian22a] 0.743$\pm$0.083 0.067$\pm$0.044
dorsogna-10k Ours (joint, v3) 0.689$\pm$0.021 0.088$\pm$0.004
Ours (PH-only, v1) 0.680$\pm$0.025 0.090$\pm$0.005
PSK [Giusti23a] 0.641$\pm$0.006 0.100$\pm$0.003
Crocker Stacks [Xian22a] 0.258$\pm$0.316 0.294$\pm$0.405
vicsek-10k Ours (joint, v3) 0.574$\pm$0.030 0.144$\pm$0.006
Ours (PH-only, v1) 0.578$\pm$0.032 0.146$\pm$0.006
PSK [Giusti23a] 0.466$\pm$0.009 0.173$\pm$0.003
Crocker Stacks [Xian22a] 0.342$\pm$0.167 0.878$\pm$0.519
volex-10k Ours (joint, v3) 0.871$\pm$0.019 0.081$\pm$0.006
Ours (PH-only, v1) 0.869$\pm$0.018 0.082$\pm$0.007
PSK [Giusti23a] 0.509$\pm$0.003 0.190$\pm$0.003
Crocker Stacks [Xian22a] -0.359$\pm$0.842 0.177$\pm$0.074
Remark 2.
A closer look at the volex-10k dataset, in particular, its governing equations in Fig. 3, shows that the cardinality of the point clouds may change over time due to cell division or death. While this is no practical limitation for our approach, our stability arguments from Eq. 3 no longer apply. We conjecture that one may be able to extend the result of [Skraba23a] to account for such cases, but this might require a case-by-case analysis and possible include additional assumptions.
5 Discussion
We introduced a framework to capture the dynamics observed in topological summary representations over time and presented one incarnation using a latent ODE on top of vectorized Vietoris-Rips persistence diagrams, addressing the problem of predicting parametrizations for models of collective behavior. Our approach successfully scales to a large number of observation sequences, requires little to no parameter tuning, and largely outperforms the current state-of-the-art, as demonstrated by several experiments on dynamic point cloud datasets with varying characteristics.
Limitation(s). One obvious limitation in the presented implementation is the reliance on Vietoris-Rips persistent homology of point clouds. In fact, the underlying simplicial complex will become prohibitively large (especially for $H_{2}$) once we scale up the number of points by, e.g., an order of magnitude. Using an Alpha [Edelsbrunner2010Computational] or a Witness complex [DeSilva04a] might be a viable alternative to mitigate this issue. Similarly, one may explore subsampling strategies for persistent homology, as in [Chazal15a], and learn a latent ODE from a combination of estimates.
Finally, we would like to reiterate that the core idea of learning the dynamics of vectorized persistence diagrams via a continuous latent variable model is not tied to one specific problem setting. In fact, one may apply our approach to other types of time-varying data (e.g., graphs) with appropriate downstream objectives. We hope that our work will inspire further research in this direction.
Societal impact. The present work mainly deals with an approach to capture the dynamics observed in data representing collective behavior. We assume that this has no direct societal impact, but point out that any application of such a model (aside from synthetic data), e.g., in the context of behavioral studies on living beings or health sciences (e.g., to study the development of cancer cells), should be carefully reviewed. This is especially advisable when drawing conclusions from missing measurements, as these are based on imputations and may be biased.
[…]
Supplementary material
[…]
Appendix A Simulation settings
As mentioned in Sec. 4.1, we create datasets of dynamic point clouds from simulations. All simulations are run with 200 points for 1,000 steps with a step size of $0.01$, and every 10th step is taken as an observation. Only for the final ablation experiment of Sec. 4.3 (i.e., Impact of observation timeframe), we simulate 20,000 time steps (of the D’Orsogna model) and extract training/testing sequences from this longer timeframe. The model parameters for these simulations are randomly sampled as specified below, and for each sampled parameter tuple, we create one simulation.
To create the dorsogna-10k dataset, we vary the following model parameters (see Fig. 3). Overall, we have four parameters that need to be predicted. As macroscopic regimes mainly depend on the ratios $C_{r}/C_{a}$ and $l_{r}/l_{a}$, cf. [Dorsogna06a, Fig. 1], we fix $C_{a}=l_{a}=1$ and sample $C_{r}=2^{t_{C}}$, $l_{r}=2^{t_{l}}$ with uniformly distributed $t_{C}\sim\mathcal{U}_{[-1,1]}$ and $t_{l}\sim\mathcal{U}_{[-1.5,0.5]}$. Similarly, we sample $\alpha=2^{t_{\alpha}}$ with $t_{\alpha}\sim\mathcal{U}_{[-2,2]}$ and $m=2^{t_{m}}$ with $t_{m}\sim\mathcal{U}_{[-2,2]}$. The model from the SiSyPHE library used to implement this simulation is AttractionRepulsion. Note, that the SiSyPHE library implements the mass $m$ in terms of the interaction radius parameter $R$, i.e., $m=R^{3}$ (as we simulate point clouds in 3D).
Remark 3.
For comparability (and interpretability of the parameters) to the original D’Orsogna model from [Dorsogna06a], we adjusted the AttractionRepulsion implementation of SiSyPHE to directly match [Dorsogna06a, Eqs. (2) & (3)].
The dorsogna-1k dataset is created by re-running the simulation provided as part of the public (Julia) implementation [^3] https://github.com/ldarrick/paths-of-persistence-diagrams of [Giusti23a]. This dataset has two parameters to predict. Note that in [Giusti23a], the authors simulated 500 sequences. For our work, we simulated 1,000 to have a larger dataset, but still one magnitude smaller than dorsogna-10k, vicsek-10k and volex-10k. For this dataset, particle masses are $m=1$, propulsion is $\alpha=1$, $C_{a}=l_{a}=1$ and $C_{r},l_{r}$ vary uniformly in $[0.1,2]$, and are selected if the generated point clouds satisfy a certain scale condition, leading to the parameter pairs illustrated in [Giusti23a, Fig. 6].
For the vicsek-10k dataset, we sample $R,c,\nu$ uniformly from $\mathcal{U}_{[0.5,5.0]}$ and $D$ from $\mathcal{U}_{[0,2.0]}$. Overall, this gives four parameters that need to be predicted.
Finally, for the volex-10k dataset, we sample four parameters as follows: $\alpha\sim\mathcal{U}_{[0,2]}$, interaction radii $R\sim\mathcal{U}_{[0,2]}$ and birth/death rates $\lambda_{b},\lambda_{d}\sim\mathcal{U}_{[0,1]}$.
However, for the latter two, we discard settings where $\lambda_{d}>\lambda_{b}$, as in this case, death rates are almost impossible to distinguish, and also settings where $\lambda_{b}\gg\lambda_{d}$ in the sense that the resulting point cloud cardinalities exceed 2000 points during the simulation. The resulting $(\lambda_{b},\lambda_{d})$ combinations are illustrated in Fig. 5.
Figure 5: Birth rates $\lambda_{b}$ and death rates $\lambda_{d}$ used for generating the volex-10k dataset.
For reproducibility, we will release the simulation data publicly.
Appendix B Comparison(s) to prior work
B.1 Path Signature Kernel (PSK) [Giusti23a]
For the PSK, we rely on the publicly available implementation in sktime444 https://github.com/sktime/sktime [Loenig19a]. We compute the PSK (truncated at signature depth 3) using our vectorized persistence diagrams per time point as input. For concatenated zero-, one- and two-dimensional persistence diagrams, using 20-dimensional vectorizations (as specified in Sec. 4.2), this yields 60-dimensional input vectors per time point. The kernel is then input to a kernel support vector regressor (kernel-SVR).
Following the experimental protocol in [Giusti23a, Sec. 7.3], we cross-validate the sliding window embedding (across lags $[1,2,3]$) and kernel-SVR hyperparameters on a 20% validation portion of the training data, with hyperparameters selected based on the average MSE (per governing equation parameter to be predicted) across 5-folds. Due to the excessive runtime (approx. 8.3 hours per lag on the system specified in Appendix C), we only report results when all time points per observation sequence are considered. Hence, no subsampling of time points (as in our approach) is performed, and the reported performance can be considered an optimistic estimate of what can be achieved with the PSK.
B.2 Crocker Stacks [Xian22a]
As mentioned in Sec. 4.4, we compare against crocker stacks, introduced in  [Xian22a], as one of the state-of-the-art approaches. Crocker stacks are an extension to crocker plots [Topaz15a] and constitute a topological summary for time-varying persistence diagrams. For the crocker stacks, we adapted the publicly available implementation of the crocker plots in teaspoon [^5] https://teaspoontda.github.io/teaspoon.
The crocker plot is computed as follows: for each time step, the persistence diagrams are computed up to a scale parameter $\varepsilon$. In discretized steps of the scale parameter, the Betti numbers are computed from the persistence diagrams, which results in a 2D representation for each homology dimension.
The extension to crocker stacks is achieved by adding a third dimension, induced by the smoothing factor $\alpha$. For given steps of $\alpha$, a smoothing operation is applied, i.e., values within a specified distance (smoothing factor) from the diagonal of the persistence diagram are ignored. Hence, for each homology dimension, a crocker stack is a tensor in $\mathbb{R}^{3}$, with axes corresponding to discretizations of the (1) scale parameter $\varepsilon\in[0,\infty)$, the (2) time $t\in[0,T]$, and the (3) smoothing factor $\alpha\in[0,\infty)$.
We note that in [Xian22a], the authors set the scale parameter to $\delta=0.35$, as they use data normalized to the range of 00 to $1$. We did not scale the data, however, we adjusted the scale parameter correspondingly: for all experiments, we set the scale parameter to $\delta=\nicefrac{{1}}{{3}}\cdot\text{maxPers}(\operatorname{dgm}_{k}(\text{Rips}(\mathcal{P})))$, where $\mathcal{P}$ denotes the point cloud and $\text{maxPers}(\operatorname{dgm}_{k}(\text{Rips}(\mathcal{P})))$ is the maximum persistence obtained from all point clouds in an observed sequence.
Hyperparameter choices. We made the following hyperparameter choices: (1) the Vietoris-Rips filtration is considered up to $\delta$ (see above), where the computations are discretized with $25$ equally spaced values in $[0,\varepsilon]$; (2) the smoothing values are discretized with $18$ steps equally spaced in $[0,0.5\cdot\text{maxPers}]$.
Eventually, the crocker stacks per homology dimension are vectorized (i.e., the tensor is flattened) and concatenated into a single vector per observation sequence. These vectorizations are then input to a linear support vector regressor (SVR). For each of the (varied) parameters in the governing equation of Fig. 3, a separate SVR is trained and evaluated. Hyperparameters of the SVR and the usage of preprocessing steps (feature scaling) are tuned using Bayesian optimization. Each parameter configuration is evaluated using a 5-fold cross-validation and the best configuration is then used to train the final model on the full dataset.
Appendix C Computational resources
All experiments were run on an Ubuntu Linux system (22.04), running kernel 5.15.0-100-generic, with 34 Intel® Core™ i9-10980XE CPU @ 3.00GHz cores, 128 GB of main memory, and two NVIDIA GeForce RTX 3090 GPUs.